OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2.

OpenStudy (shamil98):

Use the quotient rule, or turn it into the power product rule, then after you differentiate it evaluate at x = -2.

OpenStudy (anonymous):

f(x)=6x^-1 f'(x)=-6x^-2=-6/x^2 put x=-2 and get the solution.

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

First: 6/x = 6x^-1 right? Remember, exponent rules... 1/a = a^-1

OpenStudy (anonymous):

no :(

OpenStudy (agent0smith):

Well, then there it is :P 1/x = x^-1

OpenStudy (agent0smith):

So 6/x = 6x^-1 f(x) = 6x^-1 Now find f'(x) using the power rule: multiply by the exponent, then reduce the exponent by 1

OpenStudy (agent0smith):

f(x) = 6x^-1 your exponent of -1 multiplies out front, and the exponent gets reduced by 1, so -1 becomes -2: f'(x) = -1*6x^-2 f'(x) = -6x^-2 or f'(x) = -6/x^2

OpenStudy (agent0smith):

"Find the derivative of f(x) = 6/x at x = -2." now plug x=-2 into: f'(x) = -6/x^2

OpenStudy (agent0smith):

Power rule for derivatives (what I used) derivative of : x^n is nx^(n-1)